please show all work ising convolution. integral is from 0 to t

Use convolution theorem and solve y-st 0 sin(t - 2)y()dA = cost, y(0) = 1. *integral is from zero to to t I
Answer #1

Answer: The given equation is y- S* Senct-2) Y(A)da = cort - ( 1. with Y(O)=1 t Equation 0 can be written as y= Cost+ (Sin (s 2. +1 . => SL{9C4) } - {ye)ų.L {Sint} =) L{yu)} ( 3-4 s-L{ Sint} L{ych (x-1) = 1 + 1 = { county S +1 st ght an L {10} (874